OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply. (-3,4) (-1,1) A. y + 1 = -(x + 1) B. y - 4 = -(x + 3) C. y - 3 = -(x + 4) D. y - 4 = -(x + 3) E. y - 1 = -(x - 1) F. y + 1 = -(x + 1)

OpenStudy (anonymous):

picture

OpenStudy (anonymous):

the slope of the line is 1-4 / -1-(-3) = -3/2

OpenStudy (anonymous):

which means?

OpenStudy (anonymous):

hold on you got the points wrong aren't they (-3,4) and (-1-1)?

OpenStudy (anonymous):

check out the diagram - the lower point is (-1,-1)

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

agreed?

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

ok well that makes the slope (that is the gradient of the line) = to difference of the y point / difference of the x points = 4-(-1) / -3 - (-1) = 5 / -2 = -5/2 now the slope if the number whcih multiplies x in the equation of the line when we 'y' so the equation is of the form y = -5/2 x followed by some number.

OpenStudy (anonymous):

so I'm afraid it cannot be one of the eqautions you have listed. Sorry apex - I cant help you any more - something is screwed up somewhere!

OpenStudy (anonymous):

maybe we can pick the corresponding answers

OpenStudy (anonymous):

none of the answers describes the line on the diagram - just as it was on your previous post

OpenStudy (anonymous):

i can pick out any equations which fit the points - ok hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

right A and F fit (-1,-1) B,D , E fit (-3,4) C dont fit any

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok - yw plz click on good answer to close question - thanx